### v1

# Enhancing Neutrino Reconstruction in Water-Cherenkov Air Shower Arrays Using Multi-Photosensors

Abstract
In this article, the potential of water Cherenkov detectors equipped with multi-PMT modules for precision neutrino direction reconstruction is demonstrated. By analyzing signal time traces with transformer-based models, significant improvements in angular resolution are achieved compared to previous designs with larger PMTs. These detectors enable the reconstruction of neutrino directions with resolutions of approximately $10^{\circ}$ in azimuth and $7^{\circ}$ in zenith for high-signal events. This design reduces saturation effects and enhances directional sensitivity, particularly for high-energy neutrinos. The results highlight the potential of WCD arrays as complementary tools for neutrino astronomy, particularly in the context of multimessenger observations of transient astrophysical sources. The nearly continuous operation and wide field of view of these detectors further enhance their suitability for real-time monitoring and alert generation.
Keywords:
Neutrino Direction ReconstructionWater Cherenkov Detectors Multi-PMTs Multimessenger Astronomy Transformers
journal: Eur. Phys. J. C
1 Introduction
Neutrinos, being electrically neutral and weakly interacting, offer a unique window into the most extreme environments in the Universe. Unlike charged cosmic rays, which are deflected by magnetic fields, neutrinos travel freely from their sources, making them ideal messengers for identifying the origins of high-energy astrophysical phenomena. The accurate reconstruction of the neutrino direction is therefore crucial for identifying transient sources such as gamma-ray bursts, active galactic nuclei, and neutron star mergers, which are expected to produce neutrinos in the GeV to PeV energy range Ahlers:2015lln ; Ackermann:2022rqc ; Kurahashi:2022utm ; km3net2025PeV .
Large water Cherenkov detector (WCD) arrays have shown high efficacy in high-energy gamma-ray detection HAWC ; LHAASO ; SWGO . Their nearly continuous operation and wide field of view make them particularly well-suited for multimessenger astronomy Pimenta2018Astroparticle . Moreover, in previous work neutrinos_2024 , it was demonstrated that upward-going particle tracks could be identified in individual WCD stations, even when equipped with only one photosensor at the bottom and one at the top. This was achieved through the analysis of the PMT signal time traces, although with limited angular resolution. This work significantly advances those results by using multi-PMT modules, such as in various neutrino-dedicated experiments km3net_multipmt_2022 ; hyperk_multipmt_2019 , which provide enhanced directional sensitivity.
Machine learning (ML) techniques have revolutionized data analysis in astroparticle physics, particularly for ground-based observatories, where they have significantly improved performance in tasks such as event reconstruction carrillo2019improving ; GUILLEN201912 ; nieto2019ctlearn ; erdmann2018deep and gamma/hadron separation WCD_9SiPM_2020 ; alpha_2022 ; jonas_gnn_IACT_2023 ; gnn2024jonas ; ghsep_2024_tibet ; LHAASO_2020_gnn ; HAWC2021ML . Among these techniques, transformers, which are based on attention mechanisms attention_2017 , have emerged as state-of-the-art algorithms, often outperforming traditional convolutional neural networks (CNNs) footprint_2025 ; Watson2023HAWC . In this work, transformer-based models are applied to the analysis of signal time traces recorded by photomultiplier tubes (PMTs), enabling a more precise determination of neutrino direction. This approach, combined with the use of multi-PMT modules, enables the identification of smaller solid angles in the sky, improving the detection of transient astrophysical sources and complementing dedicated neutrino observatories.
The paper is structured as follows. Section 2 describes the simulation setup and detector configuration. Section 3 presents the proposed method for angular reconstruction using a single WCD equipped with two multi-PMT modules. Section 4 evaluates the performance of the method, and Section 5 presents the conclusions.
2 Detector concept and simulations
The WCD unit in this work is based on the Mercedes WCD design proposed in neutrinos_2024 ; wcd2022mercedes . The cylindrical detector retains its original dimensions (height $=1.7\,{\rm m}$, base area $\approx 12\,{\rm m^{2}}$) and features reflective interior walls to optimize Cherenkov light collection. Although the geometry remains unchanged, the photosensor configuration has been redesigned to improve the neutrino angular reconstruction performance.
In the previous design, two 8-inch PMTs were placed in the center of the upper and lower caps of the WCD (M1T1 detector). This PMT layout enabled direct detection of Cherenkov light pulses from particles entering the water volume, regardless of their direction. Although effective for neutrino detection, the azimuthal symmetry of this configuration limited its angular resolution, motivating the adoption of designs with more PMTs, such as the Mercedes WCD with three PMTs at the bottom base and one at the top (M3T1 detector).
In this work, an upgraded design is explored, replacing the two PMTs with multi-PMT modules, each comprising seven 3-inch PMTs (M1mT1m detector, see Figure 1). The PMTs are strategically arranged with a central vertical PMT and six peripherally distributed PMTs, tilted at $\theta_{\rm PMT}=45^{\circ}$ in a hexagonal pattern. With a total photosensitive area that matches the original M1T1 design, this configuration ensures at least the same detection capability while providing azimuthal sensitivity, a critical improvement for angular reconstruction. Furthermore, multi-PMT modules offer redundancy compared to larger single-PMT configurations, helping to mitigate potential saturation effects in showers with large signals.
In addition, a cost-effective alternative to the M1mT1m detector is explored, replacing the top multi-PMT module with a single 3-inch PMT (M1mT1s detector). Since Geant4 simulations are very computationally demanding, this design was implemented by reusing the simulations of the M1mT1m and excluding the six side PMTs of the upper multi-PMT module. The small PMT at the top might be particularly useful for recovering energy information in saturated stations, reducing systematic uncertainties in energy reconstruction, and improving gamma/hadron discrimination. Furthermore, its lower cost per WCD unit enables an increased array fill factor.
Figure 1: M1mT1m WCD design. The tank volume and the photocathode areas of the PMTs are represented in blue and red, respectively.
Figure 2: Mean signal time trace for nearly vertical (black histogram: $\theta_{0}\geq 175^{\circ}$, $\phi_{0}\sim 180^{\circ}$) and inclined (red histogram: $\theta_{0}\leq 150^{\circ}$, $\phi_{0}\sim 180^{\circ}$) upward-going muons with an energy of $2\,\rm GeV$. The events are centered within a radius of $r\leq 0.5\,\rm m$. The PMTs were arranged in a hexagonal layout to represent the real orientation of each PMT in the multi-PMT module.
Upward-going muon neutrinos with an energy of $1\,\rm TeV$ were simulated to evaluate the angular reconstruction capability of the detector. This energy corresponds to the intermediate energy range studied in neutrinos_2024 . The neutrinos were generated with a uniform distribution in azimuthal angle ($\phi\in[0,360^{\circ}]$) and zenith angle ($\theta\in[135^{\circ},180^{\circ}]$), and at depths uniformly distributed within $z\in[-10,0]\,\rm m$. To ensure that the neutrino interaction points are not restricted to the center of the detector, the $x$ and $y$ coordinates were shifted within a radius of 2 meters at the bottom base of the WCD. This adjustment allows the neutrinos to point to any location on the WCD bottom base.
For this study, a total of roughly 150 000150000150\,000150 000 upward-going $1\,\rm TeV$ muon-neutrinos were simulated. Only events exceeding a signal threshold of $100$ photoelectrons (p.e.) were processed, a threshold determined based on the typical signal produced by vertically crossing muons in the WCD. This cut ensures that the analysis is focused on events in which particles fully cross the detector, while effectively suppressing low-energy background sources neutrinos_2024 . After applying this cut, the dataset was divided into 44 1074410744\,10744 107 events for training, 11 0271102711\,02711 027 for validation, and 69 6526965269\,65269 652 for testing. The training set is used to optimize the model, the validation set for hyperparameter tuning and overfitting prevention, and the test set for unbiased performance evaluation.
A dedicated simulation was conducted following the approach outlined in neutrinos_2024 ; wcd2022mercedes . The detector response was simulated using Geant4 (version 4.10.05.p01) agostinelli2003geant4 ; Geant4_2016 ; WCD4PMTs , while the interactions of muon-neutrinos with the Earth’s surface were simulated using HERWIG (version herwig6521) HERWIG . The secondary particles produced in these interactions were rotated to propagate upward, aligning their trajectories toward the WCD. The subsequent development of the neutrino-induced showers was simulated in Geant4, incorporating the ground density ($\rho=2.8\,{\rm{g\,cm^{-3}}}$) and material properties, accounting for all relevant physics processes in the shower.
The analysis in this work is based on the analysis of signal time traces, which are constructed through a detailed simulation of the PMT electronics chain. For each photoelectron, a pulse is generated using reference time and amplitude distributions, and the resulting analog waveform is formed by summing these individual pulses [^1] The reference distributions for the PMT electronics simulation chain were taken from Electronics .. The waveform is then digitized by sampling its amplitude at regular intervals, determined by a sampling rate of 250 mega samples per second (MSPS). To account for potential misalignment with the master clock, the start of digitization is randomized within the time resolution. Before digitization, an offset of $0.8$ Least Significant Bit (LSB) is added to the signal. The digitization process employs a 12-bit Analog-to-Digital Converter (ADC) with a range of $[0,4095]$. In this step, the signal is truncated to the nearest integer and an offset of $400\,\rm LSB$ is added. The simulation includes PMT saturation effects, with pulses exceeding $\sim\!250\,\mathrm{p.e.}$ being clipped by the 12-bit ADC’s dynamic range.
To minimize the influence of electronics on the traces, a preprocessing step is applied after the PMT electronics simulation. First, a baseline of $402\,\rm LSB$ is subtracted from the signals to account for the offset and Gaussian noise introduced by the electronics. Second, the $T_{0}$ of the event is determined by identifying the first time bin in which the cumulative signal across all PMTs exceeds a threshold of $5\,\rm LSB$ above the baseline. This $T_{0}$ is then used to shift the time traces of all PMTs equally, aligning the event start to the beginning of the trace. In this way, the relative timing between signals in different PMTs is preserved, ensuring that temporal features such as the time differences between peaks across PMTs remain intact. Once $T_{0}$ is set, a time window of 32 ns after $T_{0}$ is used for further analysis, which is sufficient to capture the direct Cherenkov light and the first peak of reflected Cherenkov light wcd2022mercedes (see Figure 2).
The output of the simulation of the PMT electronics chain, after the signal preprocessing step, is shown in Figure 2, which shows the mean signal time traces for nearly vertical ($\theta_{0}\geq 175^{\circ}$) and inclined ($\theta_{0}\leq 150^{\circ}$) upward-going muons with an energy of $2\,\rm GeV$. The results demonstrate that this PMT layout effectively provides information about both the entry position of particles into the detector and their direction. For instance, vertical particles predominantly produce direct Cherenkov light in the top PMTs, while inclined particles can also generate direct Cherenkov light detectable in the bottom PMTs. This behavior is expected given the $41.2^{\circ}$ aperture of the Cherenkov light cone and the $45^{\circ}$ inclination of the peripheral PMTs, which optimizes their sensitivity to light from different angles.
3 Methodology for angular reconstruction
Figure 3: Flowchart of the angular reconstruction pipeline using the Transformer. Adapted from ViT_2020 .
The reconstruction of the incoming neutrino direction is performed using the signal time traces from the PMTs. Although previous work used 1-dimensional CNNs for signal time trace analysis with similar WCDs neutrinos_2024 ; wcd2022mercedes ; WCD4PMTs ; 4PMTs_NCA , the increased complexity due to the higher number of PMTs in this setup required more advanced techniques. In this work, an innovative approach is employed, using a transformer encoder attention_2017 to extract features from the signal time traces, followed by a Multi-Layer Perceptron (MLP) for regression or classification.
The transformer encoder processes the signal time traces by first splitting them into smaller segments, where each segment corresponds to the signal values of a single time-trace bin. Each segment is embedded into a higher-dimensional space, and positional encodings are added to retain temporal information. The segments are then processed through multiple transformer blocks, each consisting of self-attention and feedforward layers, to extract meaningful features. These features were subsequently used for the reconstruction of the neutrino direction.
In addition to the features extracted from the signal time traces using the transformer encoder, other engineered variables were incorporated into the MLP to enhance the reconstruction performance. These variables include: (1) total signal in the WCD; (2) the proportion of the total signal recorded by each PMT, normalized by the total signal across all PMTs; (3) the time bin corresponding to the maximum signal in each PMT; and (4) the absolute difference in time between the maximum signal in each PMT and the maximum signal in the opposite multi-PMT module. These variables, which were previously used for neutrino identification in neutrinos_2024 , provide additional contextual information on the spatial and temporal distribution of the signal.
The reconstruction process is then divided into two steps. First, given the large number of events that share the same direction but differ in vertex positions, an estimate of the entry point of the particles in the WCD is computed. This is achieved using a transformer model with a classification head at the end of the MLP. The entry points are categorized into six zones, each covering an angle of $60^{\circ}$ and centered around the position of the peripheral PMTs. For each event, the entry point $(x_{0},y_{0})$ at ground level is determined by extrapolating a straight line from the vertex point $(x_{v},y_{v},z_{v})$ along the direction of the neutrino. The true entry zone is then calculated based on the $(x_{0},y_{0})$ coordinates of the entry point.
Once the entry zone is classified, the direction of the particle is reconstructed using a transformer model with a regression head. Both the zenith angle ($\theta$) and the azimuthal angle ($\phi$) are reconstructed simultaneously by the same network, as they are intrinsically related and contribute to the observed signal time trace patterns.
The transformer models were implemented using PyTorch, with hyperparameters such as the number of layers, attention heads, and feedforward dimensions optimized using the Optuna framework optuna2019 . The optimal configuration included 2 transformer layers, 8 attention heads, a model dimension of 32, and a feedforward dimension of 128. Furthermore, the model featured two dense layers with 64 and 16 neurons, respectively, using ReLU activation and dropout rates of 0.1. The Adam optimizer kingma2014adam was used with a learning rate of $0.001$, betas of $(0.9,0.999)$, and a learning rate decay of $10^{-8}$. Training was conducted over 1 00010001\,0001 000 epochs with a batch size of 1 02410241\,0241 024 and an early stopping tolerance of 50 epochs to avoid overfitting. To ensure reproducibility and efficient monitoring of the training process, the results of the experiments were systematically tracked and managed using MLflow MLflow2018 .
For the classification of the entry zone, a standard cross-entropy loss was used. For the angular reconstruction task, on the other hand, the loss function was specifically designed to handle the periodic nature of angles, as proposed in neutrinos_2024 . A circular regression loss function was used, defined as: ${\rm Loss}=\frac{1}{2N}\sum^{N}_{i=1}\left[1-\cos(y_{i}-\hat{y}_{i})\right]\in[0,1]$. Here, $N$ is the number of samples, $y_{i}$ is the true angle for the $i$-th sample, and $\hat{y}_{i}$ is the predicted angle. This formulation ensures accurate reconstruction within the range of 0 to 360 degrees by explicitly accounting for the periodicity of angles. The loss is normalized by $1/2N$ to scale it between 0 and 1.
In summary, the angular reconstruction process was illustrated in Figure 3. The process begins with a trigger phase, during which only events exceeding the signal threshold of $100\,\rm p.e.$ are processed, determined based on the typical muon signal. Once triggered, the signal time traces for each PMT are generated through the PMT electronics simulation chain. During preprocessing, a baseline of $400\,\rm LSB$ is removed, $T_{0}$ is fixed to align the traces from different PMTs, and both the traces and engineered variables are standardized using z-score normalization. Subsequently, the particle’s entry point is classified using a transformer-based model, followed by the reconstruction of the neutrino direction using another transformer model.
4 Results
As described above, the reconstruction process begins with the estimation of the entry zone. The performance of this step is depicted in Figure 4, which shows the distribution of the angular distance in $\phi$ between the true and predicted entry zones. The plot indicates that $\sim 70\%$ of the events are reconstructed with the correct entry zone position (angular distance of $0^{\circ}$), while most of the remaining misclassified events occur near the boundaries between zones, with angular distances lower than $30^{\circ}$. This is expected, as the Cherenkov cone angle of $41.2^{\circ}$ makes it difficult to classify events near the boundaries of the $60^{\circ}$ zones, where direct light can be detected by multiple PMTs.
Figure 4: Histogram of the angular distance in $\phi$ between the true and predicted entry zones for neutrino events. An angular distance of zero denotes a correct classification of the entry zone.
It should be noted that, while the estimation of the entry zone is not essential for the final angular reconstruction, it serves as a valuable intermediate step. By providing additional context about the vertex position, the entry zone classification helps the transformer model discern patterns related to the neutrino direction, improving the overall reconstruction accuracy.
The performance of the angular reconstruction is assessed in Figure 5, which compared the efficiency of the WCD configurations using larger single PMTs, as described in previous work neutrinos_2024 , with the multi-PMT-based detectors introduced in this study. To ensure that performance differences arise solely from the detector concept and not from factors such as the reconstruction strategy or electronics, the same PMT electronics simulation and angular reconstruction strategy were applied consistently across all WCD designs. Furthermore, the hyperparameters of the transformer model were independently optimized for each WCD configuration using Optuna optuna2019 . The efficiency is quantified as the error between the true and reconstructed angles, plotted as a function of the total signal detected in the WCD. The bias of the distribution and the $68\%$ containment of the data are shown, representing the resolution of the method.
(a) Azimuthal angle.
(b) Zenith angle.
Figure 5: Comparison of the angular reconstruction efficiency as a function of the signal for the M1T1 (gray, neutrinos_2024 ), M3T1 (blue, neutrinos_2024 ), M1mT1s (dashed black line, this work), and M1mT1m (orange, this work) detectors. The top panel shows the angular reconstruction efficiency, while the bottom panel displays the number of events within each signal bin, normalized to the total number of events in the test dataset. Error bars represent the $68\%$ containment for each signal bin. Note that the M1T1 detector is not able to reconstruct $\phi$ due to the azimuthal symmetry of its PMTs.
For events with a few hundred p.e., the reconstruction errors are approximately $30^{\circ}$ and $8^{\circ}$ for the azimuthal and zenith angles, respectively. As the signal increases to thousands of p.e., the resolution improves significantly, reaching approximately $10^{\circ}$ in $\phi$ and $7^{\circ}$ in $\theta$. The difference in signal is attributed to the vertex position: for low-signal events, the vertex is typically far from the WCD, resulting in only a muon crossing the detector. In contrast, high-signal events are likely to interact closer to the detector, allowing multiple particles from the neutrino-induced shower to cross the WCD and provide additional information for reconstruction.
[…]
The resolution of the zenith angle with the M1mT1m detector improves by only a few degrees compared to detectors based on single larger PMTs. This result is expected, as the total signal in the WCD–proportional to the muon track length–along with the time difference between the maximum signals in the upper and bottom PMTs and the time trace patterns, remain robust estimators across all detector designs. However, a smaller improvement was observed for the M1mT1s configuration. This is primarily due to the reduced light collection efficiency of the single 3-inch PMT, which may miss direct Cherenkov light on the top PMT for certain particle trajectories. Furthermore, the side PMTs in the multi-PMT module are tilted by $45^{\circ}$, providing additional information about the zenith angle.
To put these results in context, dedicated neutrino experiments such as IceCube and KM3NeT/ARCA achieve angular resolutions better than $1^{\circ}$ for track-like muon neutrino events in the TeV energy range. For cascade-like events, the resolutions are $>10^{\circ}$ for IceCube and $2^{\circ}-3^{\circ}$ for KM3NeT/ARCA km3net2025PeV ; KM3NeT2019sensitivity ; icecube2013EffMass ; icecube2023observation . Although the resolution of the M1mT1m detector is less precise than dedicated experiments, it is important to note that this WCD was primarily designed for gamma-ray detection. Despite this, the achieved resolution is competitive, making the M1mT1m detector a valuable complementary tool for neutrino astronomy, particularly for high-energy events where the neutrino flux is low and every detection is critical Ackermann:2022rqc ; Kurahashi:2022utm . It should also be noted that transient sources, such as gamma-ray bursts, active galactic nuclei, and neutron star mergers, are expected to produce neutrinos in bursts waxman1997high ; ioka2005tev ; abbasi2011neutrino . In such cases, the temporal clustering of neutrino events, combined with their spatial coincidence, can provide strong evidence of their astrophysical origin, even if individual angular resolutions are limited icecube2013EffMass .
5 Conclusions
In this work, we demonstrated the potential of WCDs equipped with multi-PMT modules for neutrino direction reconstruction. The reconstruction is based on the analysis of signal time traces using transformer-based models. The multi-PMT-based WCD, featuring seven 3-inch PMTs arranged in a hexagonal pattern, provides enhanced directional sensitivity. This design enables the reconstruction of neutrino directions with resolutions of approximately $10^{\circ}$ in azimuth and $7^{\circ}$ in zenith for high-signal events, representing a substantial improvement over results obtained with WCDs using multiple larger PMTs.
The angular reconstruction strategy with this WCD benefits from the finer azimuthal angular sampling provided by the $60^{\circ}$ separation of the peripheral PMTs. Compared to alternative WCDs that use multiple larger PMTs, the multi-PMT-based WCD improves azimuthal angle resolution by a factor of two and by a few degrees for the zenith angle, reducing the bias in both cases. In addition to the enhanced angular reconstruction, multi-PMTs mitigate the saturation effects by counting with more and smaller PMTs. This helps to recover energy information in saturated stations and may improve gamma/hadron discrimination.
[…]
The achieved angular resolution, while less precise than that of dedicated neutrino experiments such as IceCube and KM3NeT/ARCA, is competitive and suitable for identifying transient neutrino sources. This makes the multi-PMT-based WCDs a valuable complementary tool for neutrino astronomy, particularly in the context of multimessenger astronomy. The ability to reconstruct neutrino directions with high precision is especially important for high-energy events, where the neutrino flux is low and every detection is critical. Furthermore, the nearly continuous operation and wide field of view of WCD arrays make them ideal for real-time monitoring and alert generation, enhancing their role in the detection of transient astrophysical sources.
Acknowledgements
We would like to thank Alberto Guillén for his valuable suggestions during the early stages of this work. This work has been financed by national funds through FCT - Fundação para a Ciência e a Tecnologia, I.P., under project PTDC/FIS-PAR/4320. B.S.G. (LIP/IST) is grateful for the financial support from the FCT PhD grant PRT/BD/151553/2021 under the IDPASC program (https://doi.org/10.54499/PRT/BD/151553/2021). J. A-M has received financial support from Ministerio de Ciencia, Innovación y Universidades/Agencia Estatal de Investigación, MICIU/AEI /10.13039/501100011033 Spain (PID2022-140510NB-I00, PCI2023-145952-2) and María de Maeztu grant CEX2023-001318-M; Xunta de Galicia, Spain (CIGUS Network of Research Centers and Consolidación 2021 GRC GI-2033 ED431C-2021/22 and 2022 ED431F-2022/15).